# Relational Observables in Group Field Theory

### v1

@@ 4.3 Perspective-reduced approach
$\hat{P}_{\chi}=\ket{0}\bra{0}+\sum_{n}\sum_{\kappa_{1},\dots,\kappa_{n}}\delta(0)^{-n}\hat{F}^{(n)}_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}(\chi)\,,$ (64)
@@
The projected space is a “synchronized Fock space”, as a generic $n$-body state |$\ket{\vec{\kappa}_{1},\chi_{1},\dots,\vec{\kappa}_{n},\chi_{n}}$ is projected onto
$\hat{P}_{\chi}\ket{\vec{\kappa}_{1},\chi_{1},\dots,\vec{\kappa}_{n},\chi_{n}}=\frac{\prod_{i=1}^{n}\delta(\chi_{i}-\chi)}{\delta(0)^{n}}\ket{\vec{\kappa}_{1},\chi,\dots,\vec{\kappa}_{n},\chi}.$ (66)
The projection $\hat{P}_{\chi}\mathcal{F}_{\text{GFT}}=\mathcal{F}_{\chi}$ thus only contains synchronized $n$-body states such as |$\ket{\vec{\kappa}_{1},\chi,\dots,\vec{\kappa}_{n},\chi}$, and can be constructed out of the field operators
@@
$[\hat{\tilde{\varphi}}_{\vec{\kappa}}(\chi),\hat{\tilde{\varphi}}^{\dagger}_{\vec{\kappa}^{\prime}}(\chi)]=\delta_{\vec{\kappa}\vec{\kappa}^{\prime}}P_{\chi}\,.$ (68)
These are the same-time commutation relations assumed in the perspective-reduced framework of [65, 66, 67, 68, 69]. Thus, $\hat{P}_{\chi}$ plays the role of a “reduction map” from $\mathcal{F}$ and $\mathcal{F}_{\chi}$. However, since $\hat{P}_{\chi}$ is a projector, this reduction map is not invertible, explicitly highlighting the distinct physical content of $\mathcal{F}$ and $\mathcal{F}_{\chi}$. In contrast, note that in the QRF context, the equivalence of the two descriptions (and thus the invertibility of the reduction map) is ensured by the removal of gauge redundancy through constraint imposition [38].
@@
where $O^{(n,m)}(\chi)$ can depend on all $\vec{g}_{i}$, $\vec{g}^{{}\,\prime}_{j}$, but all $n+m$ arguments for $\chi_{i}$ are set to $\chi$. The relation between these two definitions, as shown in A.3, is
@@
$\hat{\tilde{X}}_{\chi}\ket{\vec{\kappa}_{1},\chi,\dots,\vec{\kappa}_{n},\chi}=n\chi\ket{\vec{\kappa}_{1},\chi,\dots,\vec{\kappa}_{n},\chi}.$ (75)
@@
$\hat{\tilde{\chi}}_{\chi}(\mathop{}\!\mathrm{d}\chi)\ket{\vec{\kappa}_{1},\chi,\dots,\vec{\kappa}_{n},\chi}=\mathop{}\!\mathrm{d}\chi\,\chi\ket{\vec{\kappa}_{1},\chi,\dots,\vec{\kappa}_{n},\chi},$ (76)
@@
[…]
@@
and because $\hat{\tilde{E}}^{(P)}_{\chi}$ is hermitian, $[\hat{\tilde{E}}^{(P)}_{\chi},\hat{A}^{\dagger}]=-[\hat{\tilde{E}}^{(P)}_{\chi},\hat{A}]^{\dagger}$ for all $\hat{A}$, so
@@
~^$\hat{\tilde{\mathcal{H}}}_{\chi}\equiv\frac{1}{2}\sum_{\vec{\kappa}}\hat{\tilde{\Pi}}_{\Phi,\vec{\kappa}}(\chi)\hat{\tilde{\Pi}}_{\Phi,\vec{\kappa}}(\chi)+U[\hat{\tilde{\Phi}}]\,,$ (80)
@@
~^$\hat{\tilde{\mathcal{L}}}_{\chi}=\frac{1}{2}\sum_{\vec{\kappa}}\left(\partial_{\chi}\hat{\tilde{\Phi}}_{\vec{\kappa}}(\chi)\right)^{2}-U[\hat{\tilde{\Phi}}]\,,$ (83)
@@ Appendix A Explicit computations
A.1 Covariance
@@
$\left[\hat{E}_{\tilde{\chi}}(\mathop{}\!\mathrm{d}\tilde{\chi}),\hat{\Pi}\right]$ $=i\mathop{}\!\mathrm{d}\tilde{\chi}\sum_{n=1}^{\infty}\sum_{\vec{\kappa},\vec{\kappa}_{1},\dots\vec{\kappa}_{n}}\int\mathop{}\!\mathrm{d}\chi\left[\prod_{j=1}^{n}\mathop{}\!\mathrm{d}\chi_{j}\right]\frac{1}{n}\sum_{k=1}^{n}\delta(\chi_{k}-\tilde{\chi})$
×$\quad\times\left[\hat{\varphi}_{\vec{\kappa}}^{\dagger}(\chi)\partial_{\chi}\hat{\varphi}_{\vec{\kappa}}(\chi),F^{(n)}_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}(\chi_{1},\dots,\chi_{n})\right],$ (84)
@@
A.2 Expectation value on coherent states
@@
A.3 Equivalence between projected and sharply localized observables
@@
where $F^{(0)}\equiv\ket{0}\bra{0}$ | and where the bosonic statistics of the GFT field operator are used in the third line. Also, taking the hermitian conjugate of this equation gives
@@
The two equations (95) can be used iteratively in (71). Each integral in $\hat{O}$ over scalar field arguments is localized by the delta functions in the numerator of the above equations, and each iteration produces $[\delta(0)]^{-1/2}$; the result is

### v2

@@ 4.3 Perspective-reduced approach
$\hat{P}_{\chi}=|{0}\rangle\langle{0}|+\sum_{n}\sum_{\kappa_{1},\dots,\kappa_{n}}\delta(0)^{-n}\hat{F}^{(n)}_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}(\chi)\,$, (64)
@@
The projected space is a “synchronized Fock space”, as a generic $n$-body state $|{\vec{\kappa}_{1},\chi_{1},\dots,\vec{\kappa}_{n},\chi_{n}}\rangle$ is projected onto
$\hat{P}_{\chi}|{\vec{\kappa}_{1},\chi_{1},\dots,\vec{\kappa}_{n},\chi_{n}}\rangle=\frac{\prod_{i=1}^{n}\delta(\chi_{i}-\chi)}{\delta(0)^{n}}|{\vec{\kappa}_{1},\chi,\dots,\vec{\kappa}_{n},\chi}\rangle.$ (66)
Note, in particular, that all $n$-particle states with $\chi_{i}\neq\chi$ for any $i=1,\dots,n$, belong to the kernel of this projector. The projection $\hat{P}_{\chi}\mathcal{F}_{\text{GFT}}=\mathcal{F}_{\chi}$ thus only contains synchronized $n$-body states such as $|{\vec{\kappa}_{1},\chi,\dots,\vec{\kappa}_{n},\chi}\rangle$, and can be constructed out of the field operators
@@
$\hat{\tilde{\varphi}}_{\vec{\kappa}}(\chi),\hat{\tilde{\varphi}}^{\dagger}_{\vec{\kappa}^{\prime}}(\chi)]=\delta_{\vec{\kappa}\vec{\kappa}^{\prime}}P_{\chi}\,.$ (68)
These are the same-time commutation relations assumed in the perspective-reduced framework of [70, 71, 72, 73, 74]. Thus, $\hat{P}_{\chi}$ plays the role of a “reduction map” from $\mathcal{F}$ and $\mathcal{F}_{\chi}$. However, note that the reduction map $\hat{P}_{\chi}$ is a non-invertible projector, which explicitly highlights the distinct physical content of $\mathcal{F}$ and $\mathcal{F}_{\chi}$. In other words, $\mathcal{F}_{\chi}$ does not merely describe $\mathcal{F}$ from the perspective of $\chi$; rather, it is an entirely distinct space with different physical properties, and in particular with different commutation relations, without any requirement of compatibility with $\mathcal{F}$. In contrast, note that in the QRF context the perspective-neutral and reduced descriptions are in fact equivalent, and the reduction map is invertible, since it is a gauge redundancy that has been removed in the reduced description [37]. In this sense, the terminology “perspective-reduced” and “reduction maps” introduced here should be interpreted with appropriate caution, keeping in mind the important differences between the GFT and the QRF frameworks.
@@
where $O^{(n,m)}(\chi)$ can depend on all $\vec{g}_{i}$, $\vec{g}^{{}\,\prime}_{j}$, but all $n+m$ arguments for $\chi_{i}$ are set to $\chi$. The relation between these two definitions, as shown in A.4, is
@@
$\hat{\tilde{X}}_{\chi}|{\vec{\kappa}_{1},\chi,\dots,\vec{\kappa}_{n},\chi}\rangle=n\chi|{\vec{\kappa}_{1},\chi,\dots,\vec{\kappa}_{n},\chi}\rangle.$ (75)
@@
$\hat{\tilde{\chi}}_{\chi}(\mathop{}\!\mathrm{d}\chi)|{\vec{\kappa}_{1},\chi,\dots,\vec{\kappa}_{n},\chi}\rangle=\mathop{}\!\mathrm{d}\chi\,\chi|{\vec{\kappa}_{1},\chi,\dots,\vec{\kappa}_{n},\chi}\rangle,$ (76)
@@
$\left[\hat{\tilde{E}}^{(P)}_{\chi},\hat{\tilde{\Pi}}_{\chi}\right]=\mathop{}\!\mathrm{d}\chi\left[\hat{P}_{\chi},\hat{\tilde{\Pi}}_{\chi}\right]$ $=\frac{-i}{2\delta(0)}\mathop{}\!\mathrm{d}\chi\left[\hat{P}_{\chi},\sum_{\vec{\kappa}}\hat{P}_{\chi}\hat{\varphi}^{\dagger}_{\vec{\kappa}}(\chi)\hat{P}_{\chi}\partial_{\chi}(\hat{P}_{\chi}\hat{\varphi}_{\vec{\kappa}}(\chi)\hat{P}_{\chi})-\text{h.c.}\right];$
@@
and because $\hat{\tilde{E}}^{(P)}_{\chi}$ is hermitian, $\hat{\tilde{E}}^{(P)}_{\chi},\hat{A}^{\dagger}]=-[\hat{\tilde{E}}^{(P)}_{\chi},\hat{A}$$^{\dagger}$ for all $\hat{A}$, so
@@
$\hat{\tilde{\mathcal{H}}}_{\chi}\equiv\frac{1}{2}\sum_{\vec{\kappa}}\hat{\tilde{\Pi}}_{\Phi,\vec{\kappa}}(\chi)\hat{\tilde{\Pi}}_{\Phi,\vec{\kappa}}(\chi)+U[\hat{\tilde{\Phi}}$]\,, (80)
@@
$\hat{\tilde{\mathcal{L}}}_{\chi}=\frac{1}{2}\sum_{\vec{\kappa}}\left(\partial_{\chi}\hat{\tilde{\Phi}}_{\vec{\kappa}}(\chi)\right)^{2}-U[\hat{\tilde{\Phi}}$]\,, (83)
@@ Appendix A Explicit computations
A.1 Essentially self-adjoint
In this appendix we prove that the operator $\hat{\chi}$ given in (28) is essentially self-adjoint.
From (3.2), it is clear that the $n$-particle states are eigenvectors of $\hat{\chi}$ with real eigenvalues $\lambda_{\chi}$. As a consequence, the range of $(\hat{\chi}\pm i)$ is dense. To see this, suppose that there exists a vector $|{\psi}\rangle$ orthogonal to the range of $(\hat{\chi}\pm i)$, i.e., for any $|{\phi}\rangle\in\mathcal{F}$, $\langle{\psi}|(\hat{\chi}\pm i)|{\phi}\rangle=0$. Now take $|{\phi}\rangle$ to be an $n$-particle state, $|{n}\rangle$ (using a simplified notation here that suppresses the explicit dependence of the $n$-particle state on all the quantum numbers). Then the above condition becomes $(\lambda_{\chi}\pm i)\langle{\psi}|{n}\rangle=0$, for all $n$-particle states $|{n}\rangle$ (including the vacuum state $|{0}\rangle$ for which $\lambda_{\chi}=0$). But since $\lambda_{\chi}\in\mathbb{R}$ and the $n$-particle states form a basis for $\mathcal{F}$, we conclude that the above condition implies $|{\psi}\rangle=0$. Thus, the range of $(\hat{\chi}\pm i)$ is dense, and it follows that the operator $\hat{\chi}$ is essentially self-adjoint: it has a unique self-adjoint extension.
A.2 Covariance
@@
[…]
$\quad\times\left[\hat{\varphi}_{\vec{\kappa}}^{\dagger}(\chi)\partial_{\chi}\hat{\varphi}_{\vec{\kappa}}(\chi),F^{(n)}_{\vec{\kappa}_{1},\dots,\vec{\kappa}_{n}}(\chi_{1},\dots,\chi_{n})\right],$ (84)
@@
A.3 Expectation value on coherent states
@@
A.4 Equivalence between projected and sharply localized observables
@@
where $F^{(0)}\equiv|{0}\rangle\langle{0}|$ and where the bosonic statistics of the GFT field operator are used in the third line. Also, taking the hermitian conjugate of this equation gives
@@
The two equations (95) can be used iteratively in (71). Each integral in $\hat{O}$ over scalar field arguments is localized by the delta functions in the numerator of the above equations, and each iteration produces $\delta(0)]^{-1/2}$; the result is